a container has 8 blue balls, 6 red balls, and 4 green balls. If 3 balls are drawn, with replacement, what is the probability that all three will be green?
1
Expert's answer
2012-08-17T09:13:34-0400
P = P1 * P2 * P3
If replacement occurs, P1 = P2 = P3 = 4 / (8 + 6 + 4) = 4 /
18.
P = 4^3 / 18^3 = 64 / 5832 = 8 / 729.